# CP Violation of Baryon Decays from $N\pi$ Scatterings

### v1

@@ Abstract
Conclusions.— We propose a new CPV mechanism of baryon decays which stems from $N\pi$ scatterings. The weak phases come from the tree and penguin operators, while the strong phases are from the $N\pi$ scatterings. The most important advantage is that the fruitful data of $N\pi$ scatterings with precise partial-wave amplitudes could help us do model-independent analysis on the strong phases of $\Lambda_{b}$ decays. This is beneficial for our dynamical predictions on CPV of baryon decays. We study the CPV of $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$, $(p\pi^{0})K^{-}$, $(p\pi^{+}\pi^{-})\pi^{-}$ and $(p\pi^{+}\pi^{-})K^{-}$, via $N\pi\to p\pi^{0}$ and $N\pi\to\Delta^{++}\pi^{-}\to p\pi^{+}\pi^{-}$, respectively. It is found that the global CPV of these processes are at the order of several percents. The regional CPV of Dalitz plots could reach as large as $20\%$. The CPV of the moment $\cos\theta$ could be even larger. Such remarkable CPV have a large possibility to be observed in experiments. We strongly suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ via either the global CPV or regional CPV or CPV of moment of $\cos\theta$ to achieve the first observation of baryon CPV.
@@
Acknowledgement.—The authors are very grateful to Yanxi Zhang for the fruitful discussions on the experimental measurements at LHCb, and also to Xiaorui Lv, Wenbin Qian and Yangheng Zheng for the discussions on the prospects of experimental measurements. The authors are also grateful to Dan Guo and Igor Strakovsky for the introductions on the data of $N\pi$ scatterings from the SAID program SAID , and to Hai-Yang Cheng for the discussions on three-body $B$-meson decays. This work is supported in part by Natural Science Foundation of China under grant No. 12335003, and by the Fundamental Research Funds for the Central Universities under No. lzujbky-2024-oy02.

### v3

@@ IV Conclusions
We propose a new CPV mechanism of baryon decays which stems from $N\pi$ scatterings. The weak phases come from the tree and penguin operators, while the strong phases are from the $N\pi$ scatterings. The most important advantage is that the fruitful data of $N\pi$ scatterings with precise partial-wave amplitudes could help us do model-independent analysis on the strong phases of $\Lambda_{b}$ decays. This is beneficial for our dynamical predictions on CPV of baryon decays. We study the CPV of $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$, $(p\pi^{0})K^{-}$, $(p\pi^{+}\pi^{-})\pi^{-}$ and $(p\pi^{+}\pi^{-})K^{-}$, via $N\pi\to p\pi^{0}$ and $N\pi\to\Delta^{++}\pi^{-}\to p\pi^{+}\pi^{-}$, respectively. It is found that the global CPV of these processes are at the order of several percents. The regional CPV of Dalitz plots could reach as large as $10\%$, or be even larger. The CPV of the Legendre moments could be large as well. Such remarkable CPV have a large possibility to be observed in experiments. We strongly suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ to achieve the first observation of baryon CPV.
Acknowledgements.
@@
The authors are very grateful to Yanxi Zhang for the fruitful discussions on the experimental measurements at LHCb, and also to Jinlin Fu, Xiaorui Lü, Wenbin Qian and Yangheng Zheng for the discussions on the prospects of experimental measurements. The authors are also grateful to Dan Guo and Igor Strakovsky for the introductions on the data of $N\pi$ scatterings from the SAID program SAID , to Hai-Yang Cheng for the discussions on three-body $B$-meson decays, to Shan Cheng and Yu-Ming Wang for the discussions on the $\Lambda_{b}\to N\pi$ transition form factors, and to Cai-Dian Lü for the discussions on the CPV from final-state interactions. This work is supported in part by Natural Science Foundation of China under grant No. 12335003, and by the Fundamental Research Funds for the Central Universities under No. lzujbky-2024-oy02 and lzujbky-2023-it12.
Appendix A Data of $N\pi$ scatterings
The partial-wave amplitudes of $N\pi$ scatterings are shown explicitly for the convenience of readers. As an example, it is shown for the partial-wave amplitudes of $N\pi\to N\pi$ scatterings, $S_{11}$, $P_{11}$, $P_{13}$ and $D_{13}$. They are taken from the SAID program SAID . See Fig. 5.
Figure 5: The partial-wave amplitudes of $N\pi\to N\pi$ scatterings, $S_{11}$, $P_{11}$, $P_{13}$ and $D_{13}$. These figures are all taken from the SAID program SAID .